If CN- is a strong field ligand, rank the following coordination compounds in order of decreasing number of unpaired d-electrons: [Fe(CN6)]3- [Pd(CN6)]2- [Mo(CN6)]4-

Group of answer choices

[Fe(CN6)]3- > [Mo(CN6)]4- > [Pd(CN6)]2-

[Mo(CN6)]4- > [Fe(CN6)]3- > [Pd(CN6)]2-

[Pd(CN6)]2- > [Fe(CN6)]3- > [Mo(CN6)]4-

[Fe(CN6)]3- > [Pd(CN6)]2- > [Mo(CN6)]4-

[Mo(CN6)]4- > [Pd(CN6)]2- > [Fe(CN6)]3-
Answer #1

CNG is strong ligand ③ [Fe(CN) -3 = a + 6(-1) = -3 n=-3+6 = +3 n- +3 onidation State. Fe=[Ar] 3 do us? F e +3 =[Ar] 3dy So pa3 [Mo (CN)7-4 n+ 6(-1) =-4 n - 6 - - 4 n= - 4+6 - +2 - +2 a so Mot2 onidation State mo - [ko] 4d5 551 motz [ka] udy mot2 [ko]